Two trains left a station at the same time. One travelled north at a certain speed and th eother travelled south at twice the speed. After 4 hours, the trains were 600 miles apart. How fast was each train moving?
Answer by ankor@dixie-net.com(22740)

Two trains left a station at the same time. One traveled north at a certain speed and the other traveled south at twice the speed. After 4 hours, the trains were 600 miles apart. How fast was each train moving?
:
Let s = northbound train
Then
2s = southbound train
:
Distance = time * speed
4s + 4(2s) = 600
:
4s + 8s = 600
:
12s = 600
:
s = 600/12
:
s = 50 mph is the northbound train
Then
2(50) = 100 mph is the southbound train
:
:
Check:
4(50) + 4(100) = 600